OpenStudy (anonymous):

Three noncollinear points determine a triangle. How many triangles can be formed with 8 points, no three of which are collinear? A. 56 B. 24 C. 336 D. 6720

OpenStudy (helder_edwin):

u have to form subsets of three points out of the 8 given so u have \[ \large \binom{8}{3} \] triangles

OpenStudy (anonymous):

Yup. Exactly as ^ said. \[nCr = \frac{n!}{(n - r)!r!}\]\[8C3 = \frac{8!}{(8 - 3)!3!}\]Can you finish it up?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

harder

OpenStudy (anonymous):

Well, the following is as far as I can get you... \[8C3 = \frac{\cancel{8!}}{\cancel{5!}3!} \implies \frac{8 \times 7 \times 6 }{3 \times 2 } \implies?\]

OpenStudy (anonymous):

56

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np :)

hero (hero):

This was posted an hour ago but I'm just now receiving a message.

OpenStudy (anonymous):

lol